The length of a rectangle is 3m longer than its width.
If the perimeter of the rectangle is 34m , find its area.

Answer by simranb(63)

Hi,Hope you find your answer,
Let the width be x
then the length would be x+3
Then Perimeter=2(l+b)=34
2(x+x+3)=34
2(2x+3)=34
4x+6=34
4x=34-6
4x=28
x=28/4
x=7
Then the length =x+3
=7+3=10m
Width = x =7m
Area = l * b
= 10 * 7
= 70sqm
Cheers!
Simran